OpenStudy (anonymous):

Find the time required for a $1000 investment to (a) double at the interest intreats rate of 0.085, compounded continuously, and (b) triple at interest rate 0.085, compounded continuously.

OpenStudy (amistre64):

ln(2) and ln(3) are in your results

OpenStudy (amistre64):

time double = ln(2)/r time triple = ln(3)/r

OpenStudy (amistre64):

\[3P = Pe^{rt}\] \[\frac{3P}{P} = e^{rt}\] \[3 = e^{rt}\] \[ln(3) = ln(e^{rt})\] \[ln(3) = rt\] \[\frac{ln(3)}{r} = t\]

OpenStudy (amistre64):

same set up for the doubler

OpenStudy (anonymous):

in case of doubling it is 8.47 years

OpenStudy (anonymous):

for tripling it is 13.41 years